OpenStudy (anonymous):

Calculate each of the following limits: 1.) lim(x^2+6x-5) x->2 2.) lim(x^2-16)/(x+4) x->-4

OpenStudy (anonymous):

factor, cancel, replace \(x\) by \(-4\) for the second one

OpenStudy (anonymous):

first one plug in the number, see what you get

OpenStudy (anonymous):

i.e. compute \[2^2+6\times 2-5\]

OpenStudy (anonymous):

So then the first one would 11

OpenStudy (anonymous):

The first one can be done with plain substituion, i.e. just plug in 2 and evaluate. The second one requires some factoring and cancellation like this:\[\bf \lim_{x \rightarrow -4}\frac{ x^2-16 }{ x+4 }= \lim_{x \rightarrow -4}\frac{ (x-4)\cancel{(x+4)} }{\cancel{ x+4} }\lim_{x \rightarrow -4}= x-4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The equal sign before x-4 isn't supposed to be there lol.

OpenStudy (anonymous):

And substitute once again.

OpenStudy (anonymous):

okay. I see it, so its...simple, once again.